### v1

# The Kinematic Properties of TŻO Candidate HV 11417 with Gaia DR3

Abstract
HV 11417 is a candidate Thorne-Żytkow Object, a red supergiant with a neutron star core, located within the Small Magellanic Cloud (SMC). Previous studies have questioned, using Gaia DR2 data, whether HV 11417 was truly located at the distance of the SMC or was instead a foreground star. However, the proper motion measurement uncertainties for HV 11417 in DR2 were high. In this work, we use Gaia DR3 data to show that HV 11417 is very likely to be a true member of the SMC. We further analyze the kinematics of HV 11417 relative to its local environment, and compare it to populations of massive and evolved stars in the SMC. We find HV 11417 has a local transverse velocity of km/s, and thus qualifies as a runaway star (v 30 km/s). This runaway classification does not conclusively prove its nature as a TŻO, particularly given results from recent TŻO models, but does indicate that HV 11417 experienced a kinematic disruption in its evolution.
1 Introduction
Thorne-Żytkow Objects (TŻOs), first theorized by Thorne & Żytkow (1975, 1977), are a theoretical and intriguing avenue of binary evolution, wherein a neutron star (NS) is engulfed by the envelope of a companion red supergiant (RSG) and merges with its core, forming a single object. Significant work has gone into formalizing the physics of their formation (Taam et al., 1978; Ray et al., 1987; Podsiadlowski et al., 1995; Fryer et al., 1996; Everson et al., 2024; Hutchinson-Smith et al., 2024; Nathaniel et al., 2025; Williams et al., 2025), evolution (Cannon et al., 1992; Cannon, 1993; Biehle, 1994; Podsiadlowski et al., 1995; Farmer et al., 2023), final fates (Podsiadlowski et al., 1995; Liu et al., 2015; Moriya, 2018; Moriya & Blinnikov, 2021), and observational properties – see O’Grady et al. (2026) for a review.
To date, the best studied TŻO candidate is HV 2112 (Levesque et al., 2014), a cool, luminous star in the Small Magellanic Cloud (SMC), though its true nature has been debated (Tout et al., 2014; Maccarone & de Mink, 2016; Beasor et al., 2018; McMillan & Church, 2018; O’Grady et al., 2020, 2023; Farmer et al., 2023). A second TŻO candidate, HV 11417, also located in the SMC, was identified as a candidate by Beasor et al. (2018). The candidacy of this star has also been debated, both from the perspective of new TŻO models (Farmer et al., 2023), but also because its membership within the SMC was brought into question by O’Grady et al. (2020). Using Gaia DR2 (Gaia Collaboration et al., 2018a) kinematics, O’Grady et al. (2020) found that the proper motion of HV 11417 was marginally inconsistent with the proper motions of SMC stars, suggesting it was instead a foreground star and thus far too dim to be a TŻO. However, measurement uncertainties within Gaia DR2, especially for the of HV 11417 (), were very large.
The 3rd Data Release of Gaia (Gaia Collaboration et al., 2023) has significantly reduced proper motion measurement uncertainties, allowing for a re-analysis of the proposed location and overall kinematic properties of HV 11417. A kinematic analysis may help constrain the true identity of HV 11417, since TŻO formation necessitates a supernova explosion, which may impart a kick to the system.
In this paper, we use Gaia DR3 to analyze the kinematic properties of HV 11417 in the context of a possible TŻO identity and compared to massive stars in the SMC in general. In §2 we determine whether or not HV 11417 is located within the SMC. Then, in §3 we compare the motion of HV 11417 to other stars in its local environment, and to other populations of massive and evolved stars in the SMC. Finally, we discuss our results in §4 and conclude in §5.
2 Confirming SMC Membership
We adopt a distance of 62.44 kpc (Graczyk et al., 2020) and mean proper motions of , (Gaia Collaboration et al., 2021) for the SMC. To determine if the proper motions of HV 11417 from Gaia DR3 are consistent with membership in the SMC, we apply the method previously used by O’Grady et al. (2020, 2023, 2024), modeled after a procedure described by Gaia Collaboration et al. (2018b), for removing likely foreground sources. A given star is considered a likely foreground source if i) its parallax over parallax error , and/or ii) its proper motions fall outside the 99.5 percentile range of the proper motion phase space of a covariance matrix describing the proper motion distribution of 1,000,000 highly probable SMC members (described in detail in §2.5 of O’Grady et al. 2020). Specifically, a source is a likely foreground star if,
where is the proper motion of the star, and and are the covariance matrix (including measurement uncertainties) and median proper motions, respectively, of the highly probable SMC stars.
The proper motion values of HV 11417, shown in Table 1, are consistent with membership in the SMC ( = 0.57 10.6), and its parallax over parallax error is -2.3, well below 4. For reference, the Gaia DR2 proper motion values for HV 11417 were and . The extreme value led to the previous uncertainty in the true location of HV 11417.
| Star | RA | Dec | 111See Equation 1 | (Sig)222Astrometric Excess Noise (Significance) | RUWE | CDF333Percentage denotes how much of the cumulative density function of local stars is interior to the star | Speed444Relative to the mean speed of local stars | 555Standard deviation of the proper motion of local stars | ||
| degree | degree | (mas yr-1) | (mas yr-1) | km/s | km/s | |||||
| HV 11417 | 15.20070 | 72.85056 | 0.590.05 | 1.280.05 | 0.57 | 0.44 (34) | 1.01 | 53.5% | 5215 | 38 |
| HV 2112 | 17.51607 | 72.61461 | 1.040.03 | 1.270.03 | 2.36 | 0.24 (44) | 1.14 | 56.6% | 549 | 30 |
| M2002 1631 | 10.55134 | 73.38681 | 0.510.03 | 1.160.03 | 1.32 | 0.09 (2) | 0.98 | 33.6% | 3316 | 39 |
| M2002 15271 | 12.50778 | 72.19072 | 0.450.03 | 1.260.03 | 1.78 | 0.22 (30) | 0.97 | 75.0% | 6714 | 35 |
| M2002 30492 | 13.63938 | 73.6837 | 0.480.03 | 1.290.02 | 1.82 | 0.08 (2) | 0.95 | 61.6% | 6013 | 40 |
| M2002 37502 | 14.14036 | 72.21111 | 0.310.03 | 1.330.03 | 5.19 | 0.24 (20) | 1.58 | 95.0% | 12514 | 43 |
3 Kinematic Properties Relative to Local Environment
3.1 Local proper motion & transverse velocity of HV 11417
Here we compare the kinematics of HV 11417 relative to other stars in its local environment. We again follow the method of O’Grady et al. (2023) to analyze the kinematics of HV 11417 compared to its local environment. We query Gaia DR3 for all sources within 5′ of HV 11417, which corresponds to a radius of pc at the distance of the SMC. Foreground stars are filtered out using the same method as §2, and we additionally add a magnitude restriction of mag, following both O’Grady et al. (2023) and Phillips et al. (2024).
In Figure 1, the proper motions of HV 11417 (gold circle) and the surrounding stars (blue circles) are plotted in the top left panel (a). The white cross indicates the mean proper motion of the local stars, weighted by measurement uncertainties and excluding HV 11417, so that if HV 11417’s motion is peculiar relative to its local environment it will not bias the mean. Black lines indicate the 1-, 2-, and 3- contours of the local density distribution. HV 11417 falls just outside the 1- contour.
We further investigate the total tangential proper motion and transverse velocity of HV 11417 relative to its local environment in Figure 2, in the leftmost bar. HV 11417 has a transverse velocity, relative to nearby stars, of 5215 km/s. This is faster than 54% of the surrounding stars, and is consistent with being a runaway star (under the criterion of v 30 km/s; Blaauw 1961; Lamb et al. 2016).
3.2 Comparison to other stars
We now compare the kinematic properties of HV 11417 to the TŻO candidate HV 2112, and to populations of massive and evolved stars in the SMC.
HV 2112: The best studied TŻO candidate, the kinematic properties of HV 2112 were previously analyzed by O’Grady et al. (2023). We show HV 2112 in panel (b) of Figure 1 and as the second bar in Figure 2. HV 2112 displays very similar properties to HV 11417 – both are located near the 1- boundary of the 2D proper motion distribution of their local environments, both have transverse velocities of 50 km/s, and for both that velocity is faster than of nearby stars.
Runaway OB stars: To identify appropriate comparison stars we use The Runaways and Isolated O-Type Star Spectroscopic Survey of the SMC (RIOTS4; Lamb et al. 2016; Oey et al. 2018; Dorigo Jones et al. 2020; Phillips et al. 2024), which found that of field OB stars in the SMC are runaways (with velocities 30 km/s) and that the dynamical ejection scenario (DES, i.e. from the star’s birth cluster) dominates over the binary supernova scenario (BSS, induced by the natal kick of the explosion). We chose 4 stars from Table 1 of Phillips et al. (2024) to plot as comparative examples – M2002 1631, 15271, 30492, and 37502 (see Table 1). These stars were selected according to the following criteria: i) We consider only stars within the ‘main body’ of the SMC, not within the SMC Wing, as our covariance matrix for removing likely foreground sources is based on likely SMC stars, and the Wing has substantially different motions (Oey et al., 2018); ii) We consider only stars identified as emission line and/or high mass X-ray binaries in RIOTS4, as these classes are most likely to have experienced a BSS kick rather than a DES ejection (Dallas et al., 2022; Phillips et al., 2024); and iii) We chose stars at a range of transverse velocities corresponding to the broad range of expected TŻO velocities, plus one very fast star ( km/s, M2002 37502) that is likely the result of a combined DES+BSS ejection (Phillips et al., 2024) as an extreme example. These stars are plotted in panels (c)-(f) in Figure 1 and the rest of the bars in Figure 2. We can see that HV 11417, as well as HV 2112, indeed meet the criteria to be runaway stars, though the uncertainties for HV 11417 are large.
When comparing to the RIOTS4 sample of likely BSS kicked stars, 43% have v 30 km/s, and 15% have v 50 km/s. When compared against the complete RIOTS4 sample (of all field OB/OBe stars), 21% have v 30 km/s and 7% have v 50 km/s.
Runaway AGBs/RSGs: We compare as well to the bulk kinematic properties of RSGs, a progenitor population for TŻOs, and asymptotic giant branch (AGB) stars, the most likely stellar classification for HV 11417 if it is not a TŻO (O’Grady et al., 2023). For RSGs we use SMC sources (removing SMC Wing stars) from the catalog of Ren et al. (2021). For the AGB sample, we select oxygen-rich AGB stars (O-AGB), as HV 11417 does not demonstrate the high mass loss rate associated with carbon-rich AGB stars666From the relations of Matsuura et al. (2013) and 2MASS, Spitzer, and WISE photometry, the mass loss rate of HV 11417 is M⊙yr., in the SMC from the Spitzer-SAGE catalog (Boyer et al., 2011). For each RSG and AGB star, we remove likely foreground stars and perform the same kinematic analysis as above. We find that 33/55% of RSGs/O-AGB stars have local transverse velocities 30 km/s, 12/23% of RSGs/O-AGB stars have local transverse velocities 50 km/s, and that 13/26% of RSGs/O-AGB stars have local transverse velocities faster than 50% of other local stars.
4 Discussion
4.1 Kinematic properties of HV 11417
The improved measurement uncertainties of Gaia DR3 enable a confident placement of HV 11417 at the distance of the SMC. Its transverse velocity, relative to the mean velocities of nearby stars, is km/s, consistent with the criterion for being a runaway star ( 30 km/s). The measurement uncertainty for HV 11417 is still high (0.05 mas yr in and compared to 0.03 mas yr for HV 2112 and other comparison stars, Table 1). HV 11417 has a higher transverse velocity than of its local environment, though we also note the relatively high velocity dispersion (38 km/s) of those nearby stars.
We consider this runaway classification of HV 11417 in the context of massive and evolved star populations in the SMC. The RIOTS4 survey (Lamb et al., 2016) finds that 65% of field OB stars are runaways, and Phillips et al. (2024) find with Gaia DR3 specifically that 46% of field B stars have v 30 km/s (and 20% show v 50 km/s). Further, in this work we find that 55% of O-rich AGB stars and 33% of RSGs in the SMC have v 30 km/s (and 23% and 12% have v 50 km/s).
We note that HV 11417 has a high astrometric excess noise parameter () of 0.44 with a high significance of 34, indicating excess scatter in Gaia’s position measurements, but a Renormalized Unit Weight Error (RUWE) of 1.01, which indicates that the single-star astrometric model fits well. While photometric variability can cause elevated values (Chiavassa et al., 2022; Gandhi et al., 2022), the distance of the SMC is too large for the apparent photocenter of even a highly variable star to wobble enough to produce 0.44 mas of excess noise. Binarity can cause high values, but the excellent single-star astrometric fit (RUWE=1.01) indicates that HV 11417 is likely a single star, though the presence of a companion at a very large separation or in a face-on configuration would not be captured by the RUWE parameter (Belokurov et al., 2020). Most likely, the high and significant value of stems from the crowded nature of SMC fields (Vasiliev, 2018; Fabricius et al., 2021).
4.2 Implications for classification
If HV 11417 is a TŻO, its kinematic properties may reflect its formation history, as described in §2.1 of O’Grady et al. (2023). Leonard et al. (1994) predicts the median velocity for TŻOs formed by the natal kick launching the NS into the secondary to be km/s. Assuming generally isotropic velocities, this corresponds to a local transverse velocity of v km/s. Alternatively, for TŻOs formed from the common envelope spiral-in after a HMXB phase (Taam et al. 1978, but see also Papish et al. 2015; Hutchinson-Smith et al. 2024; Everson et al. 2024) could have velocities in the range of 10-80 km/s (van den Heuvel et al., 2000), corresponding to v 8-65 km/s.
The local transverse velocity of HV 11417 of km/s is thus consistent with either of these formation scenarios. Combined with the luminosity (log(L/L 4.92) and elemental abundance anomalies uncovered by Beasor et al. (2018) (particularly the enhanced Rb; although Beasor et al. (2018) did not detect enhanced Li, see §5.7.3 of O’Grady et al. (2023) for discussion) the properties of HV 11417 indeed align with predictions from earlier TŻO models (Thorne & Żytkow, 1975, 1977; Cannon et al., 1992; Cannon, 1993).
HV 11417 and the TŻO candidate HV 2112 share many kinematic similarities. They both have local transverse velocities greater than 50 km/s, both have velocities faster than 50% of local stars, and both lie near the 1- density contour in the 2D proper motion distribution. HV 2112 also has similar values of (0.24), significance (44), and RUWE (1.14) as HV 11417. Both are also variable stars (Soszyński et al., 2011), though HV 2112 demonstrates a much higher optical amplitude (M 4.0) than HV 11417 (M 2.4).
However, Farmer et al. (2023) recently created new models of TŻOs with MESA, finding several differences compared to the older models of Cannon et al. (1992); Cannon (1993) – though note that the structure of these new models was only computed down to the point (colloquially known as the ‘knee’) where the envelope becomes fully convective. The luminosity, metallicity, and variability period ( 1090 day fundamental mode) of HV 11417 cannot be consistently matched by any of the Farmer et al. (2023) models, thus Farmer et al. (2023) rule HV 11417 out as a TŻO candidate.
If HV 11417 is not a TŻO, its luminosity, pulsation properties, and element abundance anomalies are best explained by a massive or super-AGB star identity (Doherty et al., 2017), similar to that of HV 2112 and other similar stars explored in O’Grady et al. (2020, 2023). The runaway velocity of HV 11417 is not at odds with such a classification, as even a 5M⊙ AGB star would have a B-type main sequence progenitor, and as discussed above, nearly half of B-type main sequence stars in the SMC are runaways (Phillips et al., 2024).
5 Conclusion
In this paper we re-analyzed the kinematic properties of the TŻO candidate HV 11417 using Gaia DR3 proper motions. We confirmed that HV 11417 is highly likely to be a member of the SMC, and found that its transverse velocity, relative to its local environment, meets the criterion of a runaway star, with v km/s. This does not conclusively distinguish between possible TŻO formation channels, nor whether HV 11417 is a TŻO or some other class of cool evolved star. Given the high fraction of OB main sequence runaways (65%), and even the high fraction of runaway luminous, cool, evolved stars (33/55% for RSGs/AGBs), HV 11417 is not kinematically unusual. Its runaway velocity does indicate that it underwent some kinematically disruptive event; whether this was a dynamical ejection, an unbinding supernova kick, or a TŻO formation specifically remains uncertain. This work, along with many others, demonstrates the exceptional wealth of information within the Gaia data. We encourage further work on the kinematics of massive, evolved stars within the Magellanic Clouds, especially at the population level, enabled by Gaia data.
Acknowledgments
The author thanks Maria Drout, Katie Breivik, and Adiv Paradise for useful conversations. The author gratefully acknowledges support from the McWilliams Postdoctoral Fellowship in the McWilliams Center for Cosmology and Astrophysics at Carnegie Mellon University.
Assistance in formatting the ‘Local Stars’ element in the legend of Figure 1 and proofreading the manuscript for spelling and grammar errors was performed with Claude, Sonnet/Opus 4.5 Model, Anthropic.
Software: This work made use of the following software packages: astropy (Astropy Collaboration et al., 2013, 2018, 2022), Jupyter (Perez & Granger, 2007; Kluyver et al., 2016), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), pandas (McKinney, 2010; pandas development team, 2025), pyia (Price-Whelan & Brammer, 2021), python (Van Rossum & Drake, 2009), and scipy (Virtanen et al., 2020; Gommers et al., 2025).
This research has made use of the Astrophysics Data System, funded by NASA under Cooperative Agreement 80NSSC21M00561; the SIMBAD database and the cross-match service operated at CDS, Strasbourg, France (Wenger et al., 2000); and data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium).
Software citation information aggregated using The Software Citation Station (Wagg & Broekgaarden, 2024; Wagg et al., 2025).